OpenStudy (anonymous):

What is the estimate that results from rounding each term of thirty and one over five divided by one and ten over thirteen to the nearest whole number? 11 15 16 30

OpenStudy (anonymous):

(30(1/5) )/ (1(10/13) = x solve for x and find the nearest number in the option

OpenStudy (anonymous):

umm what

OpenStudy (anonymous):

\[(30(1/5)) \div (1(10/13)) = x\] so \[(((30*5)+1)/5) \div (((1*13) + 10)/13) = x\] \[(151/5) \div (23/13) = x\] now solve for x